OpenStudy (macabbey101):

PLEASE HELP WILL FAN AND MEDAL!!!!!!!!!!!!!!!!!!!!!!∠J and ∠M are base angles of isosceles trapezoid JKLM. If m∠J = 19x + 3, and m∠M = 13x + 15, find m∠K

OpenStudy (macabbey101):

@Mehek14

OpenStudy (macabbey101):

@K12awesomeness

OpenStudy (foxy7029):

need help

OpenStudy (foxy7029):

o :( gtg sorry

OpenStudy (macabbey101):

yes please

OpenStudy (macabbey101):

oh ok

OpenStudy (foxy7029):

;(

OpenStudy (foxy7029):

sorry

OpenStudy (welshfella):

the base angles of an isosceles trapezium are equal

OpenStudy (macabbey101):

ok. what would I do from there?

OpenStudy (welshfella):



OpenStudy (welshfella):

well since <J = < M we can write 19x + 3 = 13x + 15 solve this for x

OpenStudy (macabbey101):

so k would be 19(2)+3=41 and m would be 13(2)+15=41

OpenStudy (welshfella):

you got x = 2 right?

OpenStudy (macabbey101):

yeah

OpenStudy (welshfella):

<J = 41 not <K

OpenStudy (macabbey101):

oh sorry

OpenStudy (welshfella):

so how would you find the measure of <k using <j = 41 dgerees?

OpenStudy (macabbey101):

would you subtract 41*2 by 360 and then divide your answer by 2? would it be 139?

OpenStudy (welshfella):

yes but more simply <J + <K = 180 degrees ( because of the 2 paralell lines) So <K = 180 - 41 = 139 degrees

OpenStudy (macabbey101):

oh okay thank you so much

OpenStudy (welshfella):

yw